Required information [The following information applies to the questions displayed below.) Part 1 of 2 Davis Stores sells cloRequired Part 1 of 2 a. Use the high-low method to estimate the fixed and variable portions of store costs based on revenues.Check my work Required Part 1 of 2 a. Use the high-low method to estimate the fixed and variable portions of store costs base I'm currently stuck on this problem even after doing exactly what my book told me to do.
Answer #1

Variable cost using high low method = (Cost at highest level of revenue – cost of lowest level)/(Difference in Revenue)

= (5149-2774)/(6974-1899)

= 46.8%

Fixed cost = Total cost – variable cost

= 5149 – 6974*46.8%

= $1,885.168

b.Estimated Overhead costs = 3,300*46.8% + 1,885.168

= $3,429.568

c.Estimated costs = 18,000*46.8% + 1885.168

= $10,309.168